OpenStudy (anonymous):

A small rural town has experienced a population decrease at a rate of 25 people every 5 years since 1970. If the population of the town in 1970 was 250, which of the following equations represents the town’s population as a function of time? (Assume 1970 is represented by x = 0, 1971 is represented by x = 1, and so on). I am absolutely terrible when it comes to math!

OpenStudy (anonymous):

Forgot to add the answer choices :/ A. y = –5x + 250 B. y = –25x + 250 C. –5x + y = 250 D. –x – 25y = 250

OpenStudy (ranga):

Here y is the population and x is the year since 1970. They tell you that the "population decreases at the rate of 25 people every 5 years" slope of a line = difference in y / difference in x (or rise / run) = -25 / 5 slope m = -5 when x = 0 (that is the year 1970), the population is 250. x = 0 represents the y-intercept. So which answer choice has slope = -5 and y-intercept = 250 ?